25.0 g of an unknown metal at 90.0C is added to 50.0 g of water at 25.0 oC causing the temperature of the water rises to 29.8oC. What is the specific heat capacity and the identity of the metal?
1
Expert's answer
2012-09-14T09:34:32-0400
qmetal = qwater

By substitution, we have (metal values on the left, water values on the right):

(mass) (ΔT) (Cp) = (mass) (ΔT) (Cp)
(25.0)(90-29.8)(Cp) = (50.0)(29.8-25.0)(4.184 J g-1 °C-1)
Cp = 0.667 J g-1 °C-1

There is no metal with such Cp in tables.